A particle in a central field moves in the spiral orbit 𝑟 = 𝑐𝜃3


1
Expert's answer
2021-07-07T08:39:40-0400

Using the orbit equation


"L=mr^2\frac{d\theta}{dt}=mc^2\theta^6\frac{d\theta}{dt}"

This can be integrated to give 


"Lt = \frac{1}{7}mc^2\theta ^7"